# A Large Language Model Enhanced Sequential Recommender for Joint Video and Comment Recommendation

### v1

@@ 4.1.4. Implementation Details
4.2. Overall Results
[…]
Comment recommendation rerformance. The results of different methods on the comment recommendation task are shown in Table 3. From these results, we can find: DSSM, which does not involve user comment interaction history, exhibits inferior performance in comparison to other methods due to the lack of personalized user information. The personalized search methods that apply the title of the current video as a query (i.e., ZAM, TEM) demonstrate better performance than the simple sequential modeling methods (i.e., GRU, ATT, MHA). Based on the findings, it is evident that the personalized comment recommendation task requires user interests closely aligned with the specific video content, rather than overall generalized user comment preference. By comparing our approach with all baselines, it is clear that LSVCR achieves the most effective results.
All in all, our proposed approach shows significant improvements compared to existing baselines on both video and comment recommendation tasks. This phenomenon indicates that joint modeling of videos and comments can simultaneously enhance user video and comment preference learning. Based on personalized preference alignment, we effectively integrate the powerful semantic understanding and knowledge reasoning capabilities of the supplemental LLM recommender into the SR model. Furthermore, these results also show that the two-stage training paradigm can coordinate well for improving overall recommendation.
@@ 4.3. Ablation Study
In order to investigate how the proposed techniques affect the final performance, we conducted an ablation study on personalized video and comment recommendation tasks. Specifically, we consider the following six variants of LSVCR: (1) w/o $\mathcal{L}_{SSC}$ without the sequential-supplemental preference contrastive loss (Eq. (11)). (2) w/o $\mathcal{L}_{VCC}$ without the video-comment preference contrastive loss (Eq. (13)). (3) w/o ComDiv without the comment diversification introduced in Section 3.3.1 and maintaining inputs consistency between the two components. (4) w/o RandPosi without the randomized positional encoding introduced in Section 3.1.2. It expands the positional encoding during fine-tuning and randomly initializes the expanded part. (5) w/o Alignment without the personalized preference alignment and learning the SR model directly. (6) w/o $\mathcal{L}_{REG}$ without the ID-text regularization loss (Eq. (17)).
The results presented in Table 4 clearly demonstrate that removing any of the above techniques would lead to a decline in the overall effect. Concretely, the absence of personalized preference alignment (i.e., w/o Alignment) results in significantly weaker performance. Moreover, both contrastive losses (i.e., $\mathcal{L}_{SSC}$ and $\mathcal{L}_{VCC}$) contribute to the downstream tasks. These observations highlight the importance of our proposed various techniques in integrating the merits of the SR model and the supplemental LLM recommender.
@@ 4.4. Further Analysis
Figure 4. Performance comparison on unseen videos and users. We present the results of NDCG@10.
4.5. Online A/B Testing
In order to further verify the effectiveness of LSVCR, we deploy it to the Kuaishou platform for two weeks of online A/B testing. Specifically, we incorporate our method into the current recommendation workflow for comparison, and the experimental results are shown in Table 5. Due to the limited computational resources, we select 20K users for testing and randomly split them into two groups. We consider two metrics to measure user engagement: (1) Watch Time: Average video/comment watch time. (2) Interaction Num.: Average number of video/comment interactions. Video interactions include “like”, “collect” and “share”. Comment interactions include “like” and “reply”. The results indicate that LSVCR achieves significant improvements on both video and comment recommendation tasks. It is worth noting that the promotion of comment recommendation is much more effective compared to video recommendation which may be attributed to the integration of LLM’s powerful understanding capability on natural language in the proposed LSVCR. Overall, our framework reveals the great potential of LLM for joint video and comment recommendation on online video platforms.
Table 5. Results of online A/B testing on KuaiShou.
@@ 4.5. Online A/B Testing
5. Related Work
Sequential recommendation Sequential recommendation aims to mine users’ personalized preferences and recommend the next item for each user. This field has attracted a lot of studies due to its advantages in capturing the dynamic characteristics of user behaviors (Hidasi et al., 2016; Kang and McAuley, 2018; Sun et al., 2019). Early methods follow the Markov Chain assumption and apply matrix factorization to model item-item transfer relationships (Rendle et al., 2010; He and McAuley, 2016). Recently, typical works are mostly based on various deep neural networks to better user behavior modeling, including RNN (Hidasi et al., 2016; Li et al., 2017), CNN (Tang and Wang, 2018; Yuan et al., 2019), GNN (Wu et al., 2019a; Xu et al., 2019), and Transformer (Kang and McAuley, 2018; Sun et al., 2019). Furthermore, several studies focus on utilizing additional contextual information (e.g., title, description, category) to enhance item sequence modeling (Zhang et al., 2019; Zhou et al., 2020; Xie et al., 2022). Unlike these works, this paper specifically concentrates on video comments, incorporating abundant textual information and user feedback from comment pages into sequential recommendation.
Text-enhanced recommendation. On various application platforms, items are associated with abundant textual features, which motivates researchers to engage in text-enhanced recommendation models. Recently, pre-trained large language models (LLMs) have gained significant popularity and shown great potential in wide-ranging fields (Zhao et al., 2023; Jiang et al., 2023). In the realm of text-enhanced recommendation, LLM generally plays distinct roles: (1) LLM serves as a text feature encoder, allowing the text features to be encoded asynchronously. Subsequently, a conventional recommendation model utilizes knowledge-aware text embedding for recommendation (Wu et al., 2021; Hou et al., 2022; Yuan et al., 2023; Harte et al., 2023; Ren et al., 2023). Although effective, this approach often requires special mechanisms to enhance text embedding for better performance, such as contrastive learning (Hou et al., 2022; Ren et al., 2023), and further embedding optimization (Yuan et al., 2023; Harte et al., 2023). (2) LLM serves as a recommender. This approach typically constructs the user’s interaction history and other contexts into textual instructions, and then LLM is instructed to directly generate recommendation targets (Zhang et al., 2023b; Hou et al., 2023; Xu et al., 2024; Zheng et al., 2023; Bao et al., 2023). More recently, several works leverage various methods to enhance the adaptation of LLM for recommendation scenarios, including incorporating collaborative embedding (Zhang et al., 2023a; Li et al., 2023; Liao et al., 2023), semantic alignment (Zhu et al., 2023; Zheng et al., 2023), and so on. These methods fully utilize the excellent capabilities of LLM and achieve astonishing results, but they still cannot avoid the high computational cost and slow inference time caused by LLM. As a comparison, we aim to consolidate the merits of these two approaches through personalized preference alignment in order to achieve effective textual semantics understanding and heterogeneous interaction behavior modeling.
@@ 6. Conclusion
In this paper, we propose a novel framework called LSVCR, which leverages user interaction histories with both videos and comments for joint video and comment recommendation. There are two key components, namely conventional sequential recommendation (SR) model and supplemental LLM recommender. The former serves as the recommendation backbone to ensure efficiency, and the latter is used to enhance the semantics of the SR model. Furthermore, a two-stage training paradigm is introduced to integrate the merits of both components. For the personalized preference alignment stage, we jointly learn the video and comment recommendation tasks while performing preference alignment between the two components. For the recommendation-oriented fine-tuning stage, we discard the supplemental LLM recommender and further fine-tune the alignment-enhanced SR model to improve performance on video and comment recommendation tasks. Offline evaluation of large-scale industrial data and online A/B testing on the KuaiShou platform demonstrated the effectiveness of LSVCR. In particular, we achieve an overall watch time gain of 4.13% in the comment end. For future work, we will explore more efficient ways of utilizing LLM in joint video and comment recommendation. Additionally, we also attempt to incorporate a wider range of video and comment interaction behaviors into our framework.

### v2

@@ 5.3. Generalizability in Item Recommendation
   w/o Comment 0.3184 0.2097 - -
   w/ BERT 0.3274 0.2181 0.3851 0.2511
   w/o $L_{LM}$ 0.3206 0.2116 0.3759 0.2487
@@ 5.4. Ablation Study
In order to investigate how the proposed techniques affect the final performance, we conduct an ablation study on video and comment recommendation tasks. Specifically, we consider the following variants of LSVCR: (1) w/o $\mathcal{L}_{SSC}$ without the sequential-supplemental preference contrast (Eq. (10)). (2) w/o $\mathcal{L}_{VCC}$ without the video-comment preference contrast (Eq. (12)). (3) w/o ComDiv without the comment diversification augmentation (Section 4.3.1) and ensuring inputs consistency across the two components. (4) w/o RandPosi without the randomized positional encoding (Section 4.1.2). It expands the positional encoding during fine-tuning and randomly initializes the expanded part. (5) w/o Alignment without the personalized preference alignment and learning the SR model directly. (6) w/o $\mathcal{L}_{REG}$ without the ID-text regularization loss (Eq. (16)). (7) LSVCR w/o Comment excludes comment information and simply aligns the SR model with the LLM recommender for video recommendation task (same as the variant in Section 5.3). (8) LSVCR w/ BERT utilizes a frozen BERT as text encoder and only trains LLM recommender in the alignment stage. (9) LSVCR w/o $L_{LLM}$ without the target generation loss (Eq. (15)).
The results presented in Table 5 clearly demonstrate that removing any of the above techniques would lead to a decline in the overall effect. Concretely, the absence of personalized preference alignment (i.e., w/o Alignment) results in significantly weaker performance. Both contrastive losses (i.e., $\mathcal{L}_{SSC}$ and $\mathcal{L}_{VCC}$) contribute to the downstream tasks. Moreover, the performance degradation due to the absence of comment data once again demonstrates that comments can complement the video content, and comment interaction behaviors imply the user’s interest. It is noteworthy that our LSVCR achieves remarkable performance on both tasks, even when utilizing a frozen BERT encoder. These observations highlight the importance of our proposed various techniques in integrating the merits of the SR model and the supplemental LLM recommender.
@@ 5.5. Further Analysis
Figure 4. NDCG@10 results of unseen videos and users
@@
5.5.3. Online A/B Testing
In order to further verify the effectiveness of LSVCR, we deploy it to KuaiShou platform for 14 days of online A/B testing. Specifically, we incorporate our method into the re-ranking stage within the current recommendation workflow, and the test users are randomly divided into two groups for comparison. The experimental results are shown in Table 6. We consider two metrics to measure user engagement: (1) Watch Time: Average video/comment watch time. (2) Interaction Num.: Average number of video/comment interactions. Video interactions include “like”, “collect” and “share”. Comment interactions include “like” and “reply”. The results indicate that LSVCR achieves significant improvements on both video and comment recommendation tasks. It is worth noting that the promotion of comment recommendation is much more effective compared to video recommendation which may be attributed to the integration of LLM’s powerful understanding capability on natural language in the proposed LSVCR. Overall, our framework reveals the great potential of LLM for joint video and comment recommendation on online video platforms.
Table 6. Results of online A/B testing on KuaiShou platform.
@@ 6. Conclusion
In this paper, we proposed LSVCR that leverages user interaction histories with both videos and comments for joint video and comment recommendation. There are two key components, namely conventional sequential recommendation (SR) model and supplemental LLM recommender. The former serves as the recommendation backbone, while the latter is used to enhance the semantics of the SR model. A two-stage training paradigm is introduced to integrate the merits of both components. In the first stage, we jointly learn the video and comment recommendation tasks while performing preference alignment between two components. For the fine-tuning stage, we discard the supplemental LLM recommender and further fine-tune the alignment-enhanced SR model to improve performance on video and comment recommendation tasks. Offline evaluation and online A/B testing on a real-world platform demonstrated the effectiveness of LSVCR. As future work, we will explore more efficient ways to utilize LLM and attempt to incorporate a wider range of video and comment interaction behaviors in our framework. Additionally, inspired by the potential of multi-modal large language models (MLLMs) in cross-modal modeling, we also intend to extend LLMs to MLLMs to integrate visual features.
Acknowledgements.
This work was partially supported by National Natural Science Foundation of China under Grant No. 92470205 and 62222215, Beijing Municipal Science and Technology Project under Grant No. Z231100010323009, and Beijing Natural Science Foundation under Grant No. L233008. Xin Zhao is the corresponding author.